# Runtime Safety through Adaptive Shielding: From Hidden Parameter Inference to Provable Guarantees

### v1

@@ Abstract
Variations in hidden parameters, such as a robot’s mass distribution or friction, pose safety risks during execution. We develop a runtime shielding mechanism for reinforcement learning, building on the formalism of constrained hidden-parameter Markov decision processes. Function encoders enable real-time inference of hidden parameters from observations, allowing the shield and the underlying policy to adapt online. The shield constrains the action space by forecasting future safety risks (such as obstacle proximity) and accounts for uncertainty via conformal prediction. We prove that the proposed mechanism satisfies probabilistic safety guarantees and yields optimal policies among the set of safety-compliant policies. Experiments across diverse environments with varying hidden parameters show that our method significantly reduces safety violations and achieves strong out-of-distribution generalization, while incurring minimal runtime overhead. Project page: https://kmj1122.github.io/adaptive-shielding-project-page/
@@ 1 Introduction
Robots and other autonomous systems must operate safely in open-world environments where the underlying dynamics can vary due to hidden parameters such as mass distribution, friction, or terrain compliance. These parameters often change across episodes and remain unobserved, introducing safety risks and challenging the generalization capabilities of reinforcement learning (RL) systems (Kirk et al., 2023; Benjamins et al., 2023). Ensuring robust and safe behavior under such uncertainty is essential in domains like autonomous driving and robotic manipulation, where failures can have serious real-world consequences.
Despite recent progress in hidden parameter-aware and safe RL, existing methods often trade off adaptability and safety. Approaches such as hypernetworks, contextual models, or mixtures-of-experts (Rezaei-Shoshtari et al., 2023; Beukman et al., 2023; Celik et al., 2024) demonstrate strong adaptation to varying dynamics, but typically lack explicit mechanisms for guaranteeing safety under uncertainty. Conversely, safe RL frameworks based on constrained Markov decision processes (Wachi and Sui, 2020; Achiam et al., 2017) enforce safety by constraining cumulative costs, but often assume fixed dynamics and fail to adapt in real time to hidden parameter shifts. Methods like constrained policy optimization (Achiam et al., 2017; Yang et al., 2020) and shielding (Alshiekh et al., 2017; Yang et al., 2023) mitigate risk in static environments but remain limited in settings with dynamic or unobserved changes.
@@
To address this gap, we propose a runtime shielding framework for reinforcement learning that adapts online to hidden parameters while offering provable probabilistic safety guarantees. Central to our approach is the use of function encoders (Ingebrand et al., 2024b, 2025), a compact and expressive model class that infers environment dynamics from transition data by projecting them onto neural basis functions. This representation enables fast, online adaptation of both the policy and shield without retraining.
@@
To ensure safe learning and adaptation, our approach combines two complementary mechanisms that operate proactively during training and reactively at execution. First, we introduce a safety-regularized objective that augments rewards with a cost-sensitive value estimate, encouraging the policy to avoid unsafe behavior during training. However, this objective alone cannot guarantee safety, particularly under distribution shift. To address this, we augment policy execution with an adaptive shield that samples candidate actions from the policy, predicts their next states using a function encoder, and applies conformal prediction to quantify uncertainty in these forecasts. Actions that fail to meet a safety margin are filtered out, ensuring that only safe actions are executed.
@@
  • Online Hidden-Parameter Adaptation: We leverage function encoders to infer hidden parameters from transitions, enabling efficient policy and shield adaptation without retraining.
  • Safety-Regularized RL Objective: We propose a new objective that balances reward and safety by integrating a cost-sensitive value function, encouraging low-violation behavior during training.
  • Adaptive Shield with Probabilistic Guarantees: We develop an adaptive, uncertainty-aware runtime shield that filters unsafe actions using conformal prediction, ensuring safety during execution with provable probabilistic guarantees.
Empirical evaluations in Safe-Gym benchmarks (Ji et al., 2023), including out-of-distribution scenarios with unseen hidden parameters, demonstrate that our method significantly reduces safety violations compared to baselines, achieving robust generalization with minimal runtime overhead.
@@ 1.1 Related Work
Safe Reinforcement Learning. Safe RL methods often employ constrained MDP formulations to ensure compliance with safety constraints. Constrained policy optimization (CPO) remains foundational, effectively balancing performance and safety (Achiam et al., 2017; Wachi and Sui, 2020). Further techniques use learned recovery policy to ensure safe action execution (Thananjeyan et al., 2020). Recent zero-violation policy methods in RL aim to minimize safety violations using techniques like genetic cost function search, energy-based action filtering, and primal-dual algorithms (Hu et al., 2023; Zhao et al., 2021; Ma et al., 2024; Liu et al., 2021; Bai et al., 2023). However, these approaches often face scalability issues, rely on restrictive assumptions, or are limited to simple environments. Unlike these approaches, we introduce the safety regularized-objective that can be integrated into the optimization process of any CMDP-based RL algorithms. Shielding frameworks proactively filter unsafe actions, selectively sampling safe actions (Alshiekh et al., 2017; Carr et al., 2023; Yang et al., 2023). Recent developments on shielding integrate adaptive conformal prediction into safety frameworks, enhancing uncertainty quantification for safety-critical planning (Sheng et al., 2024a, b). However, unlike existing methods, which are not explicitly designed to address varying hidden dynamics, our approach concurrently enhances safety through a safety-regularized objective and adaptive shielding while adapting to dynamic hidden parameters using function encoders.
@@
Contextual or Hidden-Parameter Reinforcement Learning. Hidden parameters, often termed context, have been studied in recent context-aware reinforcement learning approaches, demonstrating their importance for generalization (Benjamins et al., 2023). When algorithms are provided with knowledge of the hidden parameters, they are often directly integrated into the model. For example, contextual recurrent state-space models explicitly incorporate known contextual information to enable zero-shot generalization (Prasanna et al., 2024). Contextualized constrained MDPs further integrate context-awareness into safety-prioritizing curricular learning (Koprulu et al., 2025). A common approach to handle unknown context information is to infer it from observational history using transformer models (Chen et al., 2021). Hypernetwork-based methods utilize adapter modules to adjust policy networks based on inferred contexts (Beukman et al., 2023). Mixture-of-experts architectures leverage specialized experts, using energy-based models to handle unknown contexts probabilistically (Celik et al., 2024). However, these works primarily focus on enhancing generalization to varying dynamics without incorporating safety mechanisms during adaptation in contrast to our method.
@@
Generalization in Reinforcement Learning. Generalization in RL, including zero-shot transfer and meta learning, is crucial for robust policy adaptation to varying dynamics. For example, meta-learning approaches, such as MAML (Finn et al., 2017), allow rapid parameter adaptation from minimal interaction data. Recent work in meta-safe RL (Khattar et al., 2023) has established provable guarantees for adapting to new tasks while satisfying constraints. However, meta-learning approaches involve parameter updates during adaptation, whereas our framework focuses on rapid, online inference of hidden parameters without requiring such updates. Hypernetwork-based zero-shot transfer methods explicitly condition policies on task parameters (Rezaei-Shoshtari et al., 2023). Function encoders, i.e. neural network basis functions, have demonstrated strong zero-shot transfer by using the coefficients of the basis functions as a fully-informative, linear representation of the dynamics  (Ingebrand et al., 2024b, a). Single-episode policy transfer and adaptive methods effectively handle environment changes by encoding historical context (Yang et al., 2019; Chen et al., 2022). Advanced context encoder designs further improve robustness and fast adaptation capabilities (Luo et al., 2022). While these methods excel at adapting to varying dynamics, they do not address safety constraints during adaptation, leaving agents vulnerable to unsafe actions in unseen environments.
@@ 2 Problem Formulation
Constrained hidden-parameter MDPs (CHiP-MDPs) model environments with varying transition dynamics, where a cost function is introduced alongside a reward function to address safety constraints. A CHiP-MDP extends the HiP-MDP framework Konidaris and Doshi-Velez (2014) and is defined by the tuple $\mathcal{M}=(S,A,\Theta,T,R,C,\gamma,P_{\Theta})$, where $S$ and $A$ are the state and action spaces, $R:S\times A\times S\rightarrow\mathbb{R}$ is a reward function, $C:S\times A\times S\rightarrow[0,1]$ is a cost function, and $\gamma\in(0,1)$ is the discount factor. The transition dynamics $T:S\times A\times\Theta\rightarrow S$ depend on a hidden parameter $\theta\in\Theta$. For a specified hidden parameter $\theta\in\Theta$, we denote the transition dynamics as $T_{\theta}:S\times A\rightarrow S$. The prior $P_{\Theta}(\theta)$ over the parameter space $\Theta$ represents the distribution of these hidden parameters. We denote the initial state distribution as $\mu_{0}$.
The agent follows a policy $\pi:S\times\Phi\rightarrow A$, where $\Phi$ denotes the set of learned representations of the transition dynamics $T$. Each representation $b\in\Phi$ infers the transition dynamics $T_{\theta}$ for a given parameter $\theta$. The representation $b$ can be inferred from offline or online samples from $T_{\theta}$ or replaced with $\theta$ if direct access is available. The objective of the agent is to maximize expected cumulative discounted reward while satisfying safety constraints in a CHiP-MDP $\mathcal{M}$.
To formalize this objective, we define the reward action-value function, given a parameter $\theta$, as:
$Q_{R}^{\pi}(s,a,\theta)=\mathbb{E}_{\pi,T_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},s_{t+1})\mid s_{0}=s,a_{0}=a,\theta\right].$ (1)
@@
$V_{R}^{\pi}(s,\theta)=\mathbb{E}_{a\sim\pi(\cdot\mid s,b)}\left[Q_{R}^{\pi}(s,a,\theta)\right].$ (2)
@@
$J_{R}(\pi)=\mathbb{E}_{\theta\sim P_{\Theta},s_{0}\sim\mu_{0}(\cdot\mid\theta),a_{0}\sim\pi(\cdot\mid s_{0},b)}\left[Q_{R}^{\pi}(s_{0},a_{0},\theta)\right].$ (3)
The safety constraints aim to minimize the average cost rate. To this end, we state our problem below.
Problem. Given a CHiP-MDPs $\mathcal{M}=(S,A,\Theta,T,R,C,\gamma,P_{\Theta})$ where the transition dynamics $T_{\theta}$ are fully unknown and vary with a hidden parameter $\theta$, find an optimal policy $\pi^{*}$ that maximizes the expected cumulative discounted reward $J_{R}(\pi^{*})$ while satisfying the safety constraints on the average cost rate,
$\phi^{\pi^{*}}(s,\theta)=\lim_{H\to\infty}\frac{1}{H}\mathbb{E}_{\pi^{*},T_{\theta}}\left[\sum_{t=0}^{H-1}C(s_{t},a_{t},s_{t+1})\mid s_{0}=s,\theta\right]\leq\delta,$ (4)
where $\delta\in(0,1)$ is a failure probability. Note that the transition dynamics depend on the hidden parameter $\theta$, but the policy depends on a representation of the hidden parameter derived from any previously observed transitions by $T_{\theta}$.
Lagrangian Approach. To enforce the safety constraint, we define the cost action-value function $Q_{C}^{\pi}$ and cost state-value function $V_{C}^{\pi}$ by replacing the reward function $R$ with the cost function $C$ from Equations 1 and 2. Minimizing the average cost rate can be achieved by minimizing the cost-value function $V^{\pi}_{C}$ (see Appendix I). To this end, we use a Lagrangian approach, optimizing the objective $\mathcal{L}(\pi,\lambda)=J_{R}(\pi)-\lambda(J_{C}(\pi)-\epsilon)$ with $\lambda\geq 0$, where $J_{C}(\pi)=\mathbb{E}_{\theta\sim P_{\Theta},s_{0}\sim\mu_{0}(\cdot\mid\theta)}[V_{C}^{\pi}(s_{0},\theta)]$.
@@ 3 Background
We introduce key concepts essential for understanding our methods. First, function encoders have demonstrated robust performance in estimating varying underlying dynamics (Ingebrand et al., 2024b, a, 2025). Second, conformal prediction provides a rigorous framework for quantifying uncertainty (Vovk et al., 2005; Tibshirani et al., 2019; Gibbs and Candès, 2024).
[…]
@@
For more information on how to train the neural network basis functions, see Ingebrand et al. (2025).
Conformal Prediction. Conformal Prediction (CP) allows for the construction of prediction intervals (or regions) that are guaranteed to cover the true outcome with a user-specified probability, under minimal assumptions. For exchangeable random variables $\left\{Z_{i}\right\}_{i=1}^{t+1}$, CP constructs a region satisfying: $\mathbb{P}(Z_{t+1}\leq\Gamma_{t})\geq 1-\delta,$ where $\delta\in(0,1)$ is the failure probability, and the threshold $\Gamma_{t}=Z_{(q)}$ is the $q$-th order statistic of $\left\{Z_{1},\ldots,Z_{t}\right\}$, with $q=\lceil(t+1)(1-\delta)\rceil$. Adaptive Conformal Prediction (ACP) extends this to non-stationary settings by making the threshold learnable. For more information on conformal prediction, see Shafer and Vovk (2008); Gibbs and Candès (2021).

### v2

# Adaptive Shielding for Safe Reinforcement Learning under Hidden-Parameter Dynamics Shifts

@@ Abstract
Unseen shifts in environment dynamics, driven by hidden parameters such as friction or gravity, create a challenge for maintaining safety. We address this challenge by proposing Adaptive Shielding, a framework for safe reinforcement learning in constrained hidden-parameter Markov decision processes. A function encoder infers a low-dimensional representation of the underlying dynamics online from transition data, allowing the shield to adapt. To ensure safety during this process, we use a two-layer strategy. First, we introduce safety-regularized optimization that proactively trains the policy away from high-cost regions. Second, the adaptive shielding reactively uses the inferred dynamics to forecast safety risks and applies uncertainty-aware bounds using conformal prediction to filter unsafe actions. We prove that prediction errors in the shielding connect with bounds on the average cost rate. Empirically, across Safe-Gym benchmarks with varying hidden parameters, our approach outperforms baselines on the return-safety trade-off and generalizes reliably to unseen dynamics, while incurring only modest execution-time overhead. Code is available at adaptive shield.
Safe Reinforcement Learning, Runtime Safety, Distribution Shift
@@ 1 Introduction
Robots and other autonomous systems must operate safely in real-world environments where the underlying dynamics can vary due to hidden parameters such as mass distribution, friction, or terrain compliance. These parameters often change across episodes and remain unobserved, introducing safety risks and challenging the generalization capabilities of reinforcement learning (RL) systems (kirk2023survey; benjamins2023contextualize). Ensuring robust and safe behavior under such uncertainty is essential in domains like autonomous driving and robotic manipulation, where failures can have serious real-world consequences.
Recent work has advanced along two largely separate directions. On one hand, adaptive methods such as hypernetworks, contextual policies, and mixtures-of-experts (rezaei2023hypernetworks_zero_shot; beukman2023dynamics; 2024gelik_curriculum_crl_moe) enable adaptation to changing dynamics by inferring latent environmental context, but typically do not provide explicit mechanisms for enforcing safety during this process. On the other hand, safe RL frameworks based on constrained Markov decision processes (CMDPs) (achiam2017constrained; Chen2018; wachi2020safe; yang2020PCPO; yang2022cup) enforce safety by imposing constraints on cumulative costs. Methods such as state augmentation (li2022saute) and shielding (Mohammed2017; yang2023safe) enhance safety while maintaining compatibility with a wide range of safe RL algorithms. However, these approaches often assume stationary dynamics and do not perform online adaptation under parameter shift.
@@
To address this gap, we propose an adaptive shielding framework for RL that adapts online to varying hidden parameters while offering safety bounds that are connected to the dynamics model’s prediction error. Central to our approach is the use of function encoders (tyler2024zero_shot; ingebrand2025functionencodersprincipledapproach), a compact and expressive model class that infers environment dynamics from transition data by projecting them onto neural basis functions. This low-dimensional representation enables reliable online adaptation of both the policy and shield without retraining.
@@
To ensure safe learning and adaptation, our approach combines two complementary mechanisms that operate proactively during training and reactively at execution. First, we introduce a safety-regularized optimization that augments rewards with a cost-sensitive value estimate, encouraging the policy to sample actions leading to long-term unsafe behavior during training. However, this optimization alone cannot ensure safety under distribution shift at execution. To address this, we augment policy execution with an adaptive shield that samples candidate actions from the policy, predicts future states using a function encoder, and applies conformal prediction to construct uncertainty-aware safety margins. Actions that fall below this margin are filtered out, ensuring a conservative runtime safety filter. Empirical evaluations on Safe-Gym benchmarks (ji2023safety), including out-of-distribution scenarios with unseen hidden parameters, demonstrate that our method reduces safety violations relative to baselines, while achieving robust generalization with minimal runtime overhead.
@@
  • Safety-Regularized Optimization: We introduce a training-time objective that integrates a cost-sensitive value estimate into reward maximization, optimizing the policy toward low-violation behavior.
  • Online Hidden-Parameter Adaptation: We use function encoders to infer a low-dimensional representation of hidden dynamics from transition data, enabling adaptation of both the policy and the shield without retraining.
  • Adaptive Shield with Uncertainty-Awareness: We develop an adaptive, uncertainty-aware shield that filters unsafe actions using conformal prediction, ensuring safety during execution. We analyze how prediction errors connect with the average cost rate.
@@ 1.1 Related Work
Safe Reinforcement Learning. Safe RL methods often employ constrained MDP formulations to ensure compliance with safety constraints. Constrained policy optimization (CPO) remains foundational, effectively balancing performance and safety (achiam2017constrained; wachi2020safe). Further techniques has proposed projection-based updates, recovery policies, efficient sample-manipulation, and model-based updates (zhang2020focops; yang2020PCPO; yang2022cup; gu2024enhancing; as2025actsafe). Recent zero-violation policy methods in RL aim to minimize safety violations using techniques like genetic cost function search, energy-based action filtering, primal-dual, and primal algorithms (hu2023autocost; zhao2021model; liu2021learning; bai2023achieving; ma2024learn). However, these approaches often face scalability issues, rely on restrictive assumptions, or are limited to simple environments. Unlike these approaches, we introduce the safety-regularized optimization that can be integrated into the training process of any CMDP-based RL algorithms. Shielding frameworks proactively filter unsafe actions, selectively sampling safe actions (Mohammed2017; carr2023shielding; yang2023safe). Recent developments on shielding integrate adaptive conformal prediction into safety frameworks, enhancing uncertainty quantification for safety-critical planning (lindemann2023safe_planning; sheng2024pomdp_online_shield; sheng2024daptive_conformal; scarbo2025conformal). Control barrier functions (CBFs) offer an alternative certificate-based safety mechanism. However, learning a valid barrier certificate is often difficult under uncertain or varying dynamics, as it typically requires explicit model knowledge or robust bounds on the hidden parameters (choi2020reinforcement; cheng2023safe; ganai2023iterative; wang2023enforcing; xiao2023bnet). For details on how this connects to our adaptive shielding mechanism, see Appendix D. However, unlike existing methods, which are not designed to address varying hidden dynamics, our approach concurrently enhances safety through a safety-regularized optimization and adaptive shielding while adapting to dynamic hidden parameters using function encoders.
@@
Contextual or Hidden-Parameter Reinforcement Learning. Hidden parameters, often termed context, have been studied in recent context-aware reinforcement learning approaches, demonstrating their importance for generalization (benjamins2023contextualize). When algorithms are provided with knowledge of the hidden parameters, they are often directly integrated into the model. For example, contextual recurrent state-space models explicitly incorporate known contextual information to enable zero-shot generalization (prasanna-rlc24a). Contextualized constrained MDPs further integrate context-awareness into safety-prioritizing curricular learning (koprulu2025safetyprioritizing). A common approach to handle unknown context information is to infer it from observational history using transformer models (chen2021context). Hypernetwork-based methods utilize adapter modules to adjust policy networks based on inferred contexts (beukman2023dynamics). Mixture-of-experts architectures leverage specialized experts, using energy-based models to handle unknown contexts probabilistically (2024gelik_curriculum_crl_moe). However, these works primarily focus on enhancing generalization to varying dynamics without incorporating safety mechanisms during adaptation in contrast to our method.
@@
Generalization in Reinforcement Learning. Generalization in RL, including zero-shot transfer and meta learning, is crucial for robust policy adaptation to varying dynamics. For example, meta-learning approaches, such as MAML (finn2017model), allow rapid parameter adaptation from minimal interaction data. Safe meta RL (khattar2023a; guan2024cost; xu2025efficient) extends meta-reinforcement learning to adapt to new tasks while adhering to safety constraints. However, meta-learning approaches involve parameter updates during adaptation, whereas our framework focuses on rapid, online inference of hidden parameters without requiring such updates. Hypernetwork-based zero-shot transfer methods explicitly condition policies on task parameters (rezaei2023hypernetworks_zero_shot). Function encoders, i.e. neural network basis functions, have demonstrated strong zero-shot transfer by using the coefficients of the basis functions as a fully-informative, linear representation of the dynamics  (tyler2024zero_shot; ingebrand2024zeroshot_ode). Single-episode policy transfer and adaptive methods effectively handle environment changes by encoding historical context (yang2019single; chen2022an). Advanced context encoder designs further improve robustness and fast adaptation capabilities (luo2022context). While these methods excel at adapting to varying dynamics, they do not address safety constraints during adaptation, leaving agents vulnerable to unsafe actions in unseen environments.
@@ 2 Problem Formulation
Constrained hidden-parameter MDPs (CHiP-MDPs) model environments with varying transition dynamics, where a cost function is introduced alongside a reward function to address safety constraints. A CHiP-MDP extends the HiP-MDP framework (konidaris2014hidden) and is defined by the tuple $\mathcal{M}=(S,A,\Phi,T,R,C,\gamma,P_{\Phi})$, where $S$ and $A$ are the state and action spaces, $R:S\times A\times S\rightarrow\mathbb{R}$ is a reward function, $C:S\times A\times S\rightarrow[0,1]$ is a cost function, and $\gamma\in(0,1)$ is the discount factor. The transition dynamics $T:S\times A\times\Phi\rightarrow S$ depend on a hidden parameter $\phi\in\Phi$. For a specified hidden parameter $\phi\in\Phi$, we denote the transition dynamics as $T_{\phi}:S\times A\rightarrow S$. The prior $P_{\Phi}(\phi)$ over the parameter space $\Phi$ represents the distribution of these hidden parameters. We denote the initial state distribution as $\mu_{0}$.
Since the hidden parameters $\phi$ are unknown to the agent, it must infer changes in the environment dynamics from observations. To this end, the agent follows a policy $\pi:S\times\mathcal{B}\rightarrow A$, where $\mathcal{B}$ denotes the set of learned representations of the transition dynamics $T_{\phi}$. We denote the resulting representation by $b_{\phi}$ for each $\phi$. The objective is to maximize expected cumulative discounted reward while satisfying safety constraints in a CHiP-MDP $\mathcal{M}$. To formalize this objective, we define the reward action-value function, for a parameter $\phi$, initial state $s$, and action $a$ as:
$Q_{R}^{\pi}(s,a,b_{\phi})=\mathbb{E}_{\pi,T_{\phi}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},s_{t+1})\mid s,a,\phi\right].$ (1)
@@
$V_{R}^{\pi}(s,b_{\phi})=\mathbb{E}_{a\sim\pi(\cdot\mid s,b_{\phi})}\left[Q_{R}^{\pi}(s,a,b_{\phi})\right].$ (2)
@@
$J_{R}(\pi)=\mathbb{E}_{\phi\sim P_{\phi},s_{0}\sim\mu_{0}(\cdot\mid\phi),a_{0}\sim\pi(\cdot\mid s_{0},b_{\phi})}\left[Q_{R}^{\pi}(s_{0},a_{0},b_{\phi})\right].$ (3)
Likewise, the cost objective is defined the same way, replacing the reward function $R$ with the cost function $C$: $J_{C}(\pi)=\mathbb{E}_{\phi\sim P_{\phi},s_{0}\sim\mu_{0}(\cdot\mid\phi),a_{0}\sim\pi(\cdot\mid s_{0},b_{\phi})}\left[Q_{C}^{\pi}(s_{0},a_{0},b_{\phi})\right].$ The safety constraints aim to minimize the average cost rate. To this end, we state our problem below.
Problem. Given a CHiP-MDPs $\mathcal{M}=(S,A,\Phi,T,R,C$$\gamma,P_{\Phi})$ where the transition dynamics $T_{\phi}$ are fully unknown and vary with a hidden parameter $\phi$, find an optimal policy $\pi^{*}$ that maximizes the expected cumulative discounted reward $J_{R}(\pi^{*})$ while satisfying the safety constraints on the average cost rate,
$\xi^{\pi^{*}}(s,\phi)$ $=\lim_{H\to\infty}\frac{1}{H}\mathbb{E}_{\pi^{*},T_{\phi}}\left[\sum_{t=0}^{H-1}C(s_{t},a_{t},s_{t+1})\mid s,\phi\right]$
$\leq\delta,$ (4)
where $\delta\in(0,1)$ is a failure probability and initial state is $s$. Note that the transition dynamics depend on the hidden parameter $\phi$, but the policy depends on a representation of the hidden parameter, $b_{\phi}$, derived from any previously observed transitions by $T_{\phi}$.
To enforce the safety constraint, we define the cost action-value function $Q_{C}^{\pi}$ and cost state-value function $V_{C}^{\pi}$ by replacing the reward function $R$ with the cost function $C$ from Equations 1 and 2. Minimizing the average cost rate can be achieved by minimizing the cost-value function $V^{\pi}_{C}$ (see Appendix M).
@@ 3 Background
We introduce key concepts essential for understanding our methods. First, function encoders have demonstrated robust performance in estimating varying underlying dynamics (tyler2024zero_shot; ingebrand2024zeroshot_ode; ingebrand2025functionencodersprincipledapproach). Second, conformal prediction provides a rigorous framework for quantifying uncertainty (vovk2005algorithmic; tibshirani2019conformal; gibbs2024online).
Function Encoder. A function encoder (FE) offers a compact and computationally efficient framework for representing functions in terms of neural network basis functions. Consider a set of functions $\mathcal{F}=\{f\mid f:\mathcal{X}\rightarrow\mathbb{R}\}$, where $\mathcal{X}\subset\mathbb{R}^{n}$ is an input space with finite volume. When $\mathcal{F}$ forms a Hilbert space with the inner product $\langle f,g\rangle=\int_{\mathcal{X}}f(x)g(x)dx$, any $f\in\mathcal{F}$ can be expressed using a basis $\left\{g_{1},g_{2},\ldots,g_{k}\right\}$ as $f(x)=\sum_{i=1}^{k}b_{i}g_{i}(x)$, where $b_{i}$ are unique coefficients. To determine the coefficients, we solve the following least-squares optimization problem:
@@
For more information on how to train the neural network basis functions, see ingebrand2025functionencodersprincipledapproach.
Conformal Prediction. Conformal Prediction (CP) allows for the construction of prediction intervals (or regions) that are guaranteed to cover the true outcome with a user-specified probability, under minimal assumptions. For exchangeable random variables $\left\{Z_{i}\right\}_{i=1}^{t+1}$, CP constructs a region satisfying: $\mathbb{P}(Z_{t+1}\leq\Gamma_{t})\geq 1-\delta,$ where $\delta\in(0,1)$ is the failure probability, and the threshold $\Gamma_{t}=Z_{(q)}$ is the $q$-th order statistic of $\left\{Z_{1},\ldots,Z_{t}\right\}$, with $q=\lceil(t+1)(1-\delta)\rceil$. Adaptive Conformal Prediction (ACP) extends this to non-stationary settings by making the threshold learnable. For more information on conformal prediction, see shafer2008tutorial; gibbs2021adaptive.